# Federated Aggregation of Demand Flexibility

### v1

@@ I Introduction
This critical need for privacy-preserving yet high-performing aggregation methods for demand flexibility has spurred the search for new decentralized architectures. Federated learning, in particular, offers a powerful conceptual blueprint. It enables participants to collaboratively build a shared model by exchanging only intermediate parameters, ensuring that raw data remain local, which provides a promising framework for resolving the aforementioned conflict [5].
@@
Despite these structural advantages, this “proto-federated” approach diverges from a true federated learning process. Its core limitation is that the shared model, the base set, is treated as a static, heuristic template, rather than as an optimizable variable to be collaboratively updated based on DSRs’ local data [6]. This static-template design treats DSRs as passive participants instead of active learners who can jointly update the shared model tailored to the population’s unique characteristics. This absence of a collaborative optimization process may create an unavoidable performance bottleneck for demand flexibility aggregation.
@@ II Related Literature
Existing literature on aggregating demand flexibility can be broadly classified into three categories: boundary aggregation, optimization-based and geometric methods.
@@ II-1 Boundary aggregation methods
The studies in this category model each DSR as a virtual battery by identifying the upper and lower bounds of realizable charging power and cumulative energy [9]. These individual boundaries are further aggregated to form a larger virtual battery for aggregate flexibility characterization. For example, Zhang et al. [10] and Brinkel et al. [11] first derive the realizable upper and lower power-and-energy envelopes for each individual DSR and then sum these envelopes to obtain aggregate flexibility bounds. However, this approach typically assumes homogeneous DSR behavior (e.g., similar plug-in durations) and the resulting power-and-energy envelopes may be outer approximations, in which many trajectories cannot actually be scheduled for individual DSRs [12].
@@ III-B1 Individual flexibility set
We consider a finite-horizon discrete-time model with time slots indexed by $t\in\mathcal{T}:=\{1,\dots,T\}$, and there are $N$ EVs indexed by $\mathcal{N}:=\{1,\dots,N\}$. We denote the net charging power and cumulative charging energy as $\mathbf{u}_{i}\!:=\!\big[u_{i}(t)\big]_{\!t\in\mathcal{T}}$ and $\mathbf{x}_{i}\!:=\!\big[x_{i}(t)\big]_{\!t\in\mathcal{T}}$, respectively. Given a charging profile $\mathbf{u}_{i}$, the cumulative energy stored in the EV battery can be expressed as [19]: [^1] We assume there is no energy loss during charging and discharging.
@@ III-C Revisit of Geometric Methods
s.t. $\bm{\gamma}_{i}+\bm{\Gamma}_{i}\,\mathcal{U}_{0}(\mathbf{h}_{0})\subseteq\mathcal{U}_{i},\ \ i\in\mathcal{N}.$ (8b)
@@
Note that we remove the term $\mathrm{vol}(\mathcal{U}_{0}(\mathbf{h}_{0}))$ from the objective, as it is a constant when $\mathbf{h}_{0}$ is given. Obviously, the objective (10a) couples different DSRs. As a result, directly solving (10) requires that all DSRs share their sensitive information $\mathbf{h}_{i}$ with the aggregator, which may cause privacy concerns. To address the privacy issue, reference [18] proposed to use the first-order Taylor expansion about the identity matrix to approximate the objective, as follows:
@@ III-D Base Set Optimization
Solving problem (13) is challenging, because i) it is a bilevel problem due to constraint (13c), ii) the objective function is non-concave because both $\bm{\Gamma}_{i}^{\star}$ and $\mathbf{h}_{0}$ are variables, iii) accurately computing the volume of the base set $\mathcal{U}_{0}(\mathbf{h}_{0})$ is computationally demanding, and iv) the constraint cannot be directly handled by existing optimization solvers. Together, these challenges render problem (13) intractable for off-the-shelf solvers. Moreover, this centralized formulation requires DSRs to share the vectors $\mathbf{h}_{i}$ for all $i\in\mathcal{N}$ with the aggregator, making it difficult to preserve user privacy.
@@ IV-B Federated Gradient-based Algorithm
  • Central Computation: The aggregator computes all the other terms that do not depend on private DSR data, including the volume $\phi(\widetilde{\mathbf{h}}_{0})$, its gradient $\nabla_{\widetilde{\mathbf{h}}_{0}}\phi(\widetilde{\mathbf{h}}_{0})$, and the projection gradient $\nabla_{{\mathbf{h}}_{0}}\widetilde{\mathbf{h}}_{0}({\mathbf{h}}_{0})$. Upon receiving $(\boldsymbol{\Gamma}_{i}^{\star},\nabla_{\widetilde{\mathbf{h}}_{0}}\boldsymbol{\Gamma}_{i}^{\star})$ from all DSRs, the aggregator computes $\ell(\widetilde{\mathbf{h}}_{0})$ and $\nabla_{\widetilde{\mathbf{h}}_{0}}\ell(\widetilde{\mathbf{h}}_{0})$. Crucially, since these terms involve only aggregated sums, they can be computed using secure summation protocols that preserve anonymity and prevent disclosure of user contributions. Finally, the overall gradient can be computed according to (17)-(19).
@@ IV-C1 Gradient $\nabla_{\mathbf{h}_{0}}\widetilde{\mathbf{h}}_{0}({\mathbf{h}}_{0})$
Calculating the volume of a high-dimensional polytope, such as $\phi(\mathbf{h}_{0})$, is generally intractable [23]. Existing volume approximation methods typically rely on Monte Carlo sampling [24, 25], where the volume is inferred based on the proportion of sampled points within the polytope. However, these methods are not suitable for computing gradients, as they do not provide a differentiable formulation of the volume. To address this issue, we derive analytical expressions for both the volume and its gradient. For an arbitrary base set parameter $\mathbf{h}_{0}$, the volume of the base set defined in (5) can be expressed as
@@ VI-A1 Parameter settings
We consider a finite time horizon with each time interval being 1 hour ($\Delta t=1$), and an EV fleet consisting of 50 vehicles ($N=50$). The parameter intervals of EVs are summarized in Table I (the parameters of a specific EV are uniformly sampled from these intervals). The energy demand for EV $i$ is computed by first determining its maximum feasible charge, given by $u_{i}^{\mathrm{max}}(t_{i}^{\mathrm{d}}-t_{i}^{\mathrm{p}})$, and then scaling it by a random number drawn uniformly from $[0.2,0.5]$.
@@
VI-A2 Benchmarks
@@
VI-A3 Test platform
All the case studies are implemented on a MacBook Pro 3 with 32 GB of memory. Gradients involving convex optimizations are calculated using PyTorch with differentiable convex optimization layers, CVXPYlayers [26].
@@ VI-C2 Electricity cost minimization
$\min_{\mathbf{u}_{i},\forall i\in\mathcal{N}}\ \bigg(\!\bm{\pi}^{\top}\sum_{i\in\mathcal{N}}\mathbf{u}_{i}\!\bigg)\Delta t\ \ \text{s.t. }\ \mathbf{u}_{i}\in\mathcal{U}_{i},\,\forall i\in\mathcal{N}.$ (29)
Fig. 6 illustrates the relative electricity cost gaps of the AVG and Proposed methods against the best-case benchmark, and the electricity cost improvement of the Proposed method over the AVG method. While both methods incur some level of optimality gaps compared to the best-case benchmark, our Proposed method consistently demonstrates superior performance. Employing the optimized base-set parameter $\mathbf{h}_{0}^{\mathrm{opt}}$ markedly narrows the gap. As highlighted by the third boxplot, this translates into a substantial relative cost savings, with our approach consistently outperforming the AVG benchmark by a significant margin. These results provide strong evidence that a more accurate flexibility characterization, as enabled by our method, directly translates into economic value.
@@ Appendix B Gradient Derivation for Set Containment Problem
$\nabla_{\bm{\gamma}}\mathcal{L}=\mathbf{H}^{\top}\bm{\mu}^{\star}=\mathbf{0},$ (35a)
$\nabla_{\bm{\Gamma}}\mathcal{L}=-\mathbf{I}_{T}-\mathbf{H}^{\top}\mathbf{Y}^{\star}=\mathbf{0},$ (35b)
$\nabla_{\bm{\Lambda}}\mathcal{L}=\mathbf{Y}^{\star}\mathbf{H}^{\top}+\bm{\mu}^{\star}{\mathbf{h}}_{0}^{\top}-\bm{\nu}^{\star}=\mathbf{0},$ (35c)
$\bm{\Lambda}^{\star}\geq\bm{0},\quad\bm{\Lambda}^{\star}\mathbf{H}-\mathbf{H}\bm{\Gamma}^{\star}=\mathbf{0},$ (35d)
$\mathbf{g}^{\star}=\bm{\Lambda}^{\star}{\mathbf{h}}_{0}+\mathbf{H}\bm{\gamma}^{\star}-\mathbf{h}\leq\mathbf{0},$ (35e)
$\bm{\mu}^{\star}\geq\mathbf{0},\quad\bm{\nu}^{\star}\geq\mathbf{0},\quad\bm{\mu}^{\star}\circ\mathbf{g}^{\star}=\mathbf{0},\quad\bm{\nu}^{\star}\circ\bm{\Lambda}^{\star}=\mathbf{0},$ (35f)

### v2

@@ I Introduction
This critical need for privacy-preserving yet high-performing aggregation methods for demand flexibility has spurred the search for new decentralized architectures. Federated learning, in particular, offers a powerful conceptual blueprint. It enables participants to collaboratively build a shared model by exchanging only intermediate parameters while keeping raw data local; analogously (though not for machine learning model training), this motivates a framework with similar information-exchange patterns and privacy properties for aggregation of demand flexibility to resolve the aforementioned conflict [5].
@@
Despite these structural advantages, this “proto-federated” approach diverges from a true federated process. Its core limitation is that the shared model, the base set, is treated as a static, heuristic template, rather than as an optimizable variable to be collaboratively updated based on DSRs’ local data [6]. This static-template design treats DSRs as passive participants instead of active learners who can jointly update the shared model tailored to the population’s unique characteristics. This absence of a collaborative optimization process may create an unavoidable performance bottleneck for demand flexibility aggregation.
@@ II Related Literature
Existing literature on aggregating demand flexibility [^1] Beyond explicitly aggregating the demand flexibility sets, there is a broader literature on coordinating DSRs (cf. [9, 10] and references therein). can be broadly classified into three categories: boundary aggregation, optimization-based and geometric methods.
@@ II-1 Boundary aggregation methods
The studies in this category model each DSR as a virtual battery by identifying the upper and lower bounds of realizable charging power and cumulative energy [11, 12]. These individual boundaries are further aggregated to form a larger virtual battery for aggregate flexibility characterization. For example, Zhang et al. [13] and Brinkel et al. [14] first derive the realizable upper and lower power-and-energy envelopes for each individual DSR and then sum these envelopes to obtain aggregate flexibility bounds. However, this approach typically assumes homogeneous DSR behavior (e.g., similar plug-in durations) and the resulting power-and-energy envelopes may be outer approximations, in which many trajectories cannot actually be scheduled for individual DSRs [15].
@@ III-B1 Individual flexibility set
We consider a finite-horizon discrete-time model with time slots indexed by $t\in\mathcal{T}:=\{1,\dots,T\}$, and there are $N$ EVs indexed by $\mathcal{N}:=\{1,\dots,N\}$. We denote the net charging power and cumulative charging energy as $\mathbf{u}_{i}\!:=\!\big[u_{i}(t)\big]_{\!t\in\mathcal{T}}$ and $\mathbf{x}_{i}\!:=\!\big[x_{i}(t)\big]_{\!t\in\mathcal{T}}$, respectively. Given a charging profile $\mathbf{u}_{i}$, the cumulative energy stored in the EV battery can be expressed as [22]: [^2] We adopt a lossless charging model to ensure that individual EV flexibility sets admit a convex polyhedral representation (H-representation), which is the key structural requirement of the proposed framework. The framework directly applies to one-way charging scenarios, where $u_{i}(t)\geq 0$ for all $t\in\mathcal{T}$, by incorporating charging inefficiencies as constant scaling factors in the constraint matrices. For general bidirectional lossy charging models, convex polyhedral inner approximations of the resulting nonconvex flexibility sets, admitting an H-representation, have been established in existing literature [23, 24]. Consequently, the proposed framework remains applicable when such approximations are employed.
@@ III-C Revisit of Geometric Methods
s.t. $\bm{\gamma}_{i}+\bm{\Gamma}_{i}\,\mathcal{U}_{0}(\mathbf{h}_{0})\subseteq\mathcal{U}_{i},\ \ i\in\mathcal{N},$ (8b)
@@
Note that we remove the term $\mathrm{vol}(\mathcal{U}_{0}(\mathbf{h}_{0}))$ from the objective, as it is a constant when $\mathbf{h}_{0}$ is given. Obviously, the objective (10a) couples different DSRs. As a result, directly solving (10) requires that all DSRs share their sensitive information $\mathbf{h}_{i}$ with the aggregator, which may cause privacy concerns. To address the privacy issue, we linearize the objective using a first-order Taylor expansion about the identity matrix [^3] The approximation is introduced to enable separability of the objective and enable a federated implementation. Similar linearizations of determinant-based volume metrics are commonly adopted in geometric methods and are empirically validated to provide tight aggregate flexibility representations [21]., as follows:
@@ III-D Base Set Optimization
Solving problem (13) is challenging, because i) it is a bilevel problem due to constraint (13c), ii) the objective function is non-concave because both $\bm{\Gamma}_{i}^{\star}$ and $\mathbf{h}_{0}$ are variables, iii) accurately computing the volume of the base set $\mathcal{U}_{0}(\mathbf{h}_{0})$ is computationally demanding, and iv) the constraint cannot be directly handled by existing optimization solvers. Together, these challenges render problem (13) intractable for off-the-shelf solvers. Moreover, this centralized formulation requires DSRs to share the vectors $\mathbf{h}_{i}$ for all $i\in\mathcal{N}$ with the aggregator. Since $\mathbf{h}_{i}$ encodes user preferences, schedules, and energy usage patterns, such disclosure directly exposes sensitive operational information and undermines user privacy.
@@ IV-B Federated Gradient-based Algorithm
  • Central Computation: The aggregator computes all the other terms that do not depend on private DSR data, including the volume $\phi(\widetilde{\mathbf{h}}_{0})$, its gradient $\nabla_{\widetilde{\mathbf{h}}_{0}}\phi(\widetilde{\mathbf{h}}_{0})$, and the projection gradient $\nabla_{{\mathbf{h}}_{0}}\widetilde{\mathbf{h}}_{0}({\mathbf{h}}_{0})$. Upon receiving $(\boldsymbol{\Gamma}_{i}^{\star},\nabla_{\widetilde{\mathbf{h}}_{0}}\boldsymbol{\Gamma}_{i}^{\star})$ from all DSRs, the aggregator computes $\ell(\widetilde{\mathbf{h}}_{0})$ and $\nabla_{\widetilde{\mathbf{h}}_{0}}\ell(\widetilde{\mathbf{h}}_{0})$. Crucially, since these terms involve only aggregated sums, they can be computed using secure summation protocols that preserve anonymity and substantially limit the disclosure of user contributions. Finally, the overall gradient can be computed according to (17)-(19).
@@ IV-C1 Gradient $\nabla_{\mathbf{h}_{0}}\widetilde{\mathbf{h}}_{0}({\mathbf{h}}_{0})$
Calculating the volume of a high-dimensional polytope, such as $\phi(\mathbf{h}_{0})$, is generally intractable [27]. Existing volume approximation methods typically rely on Monte Carlo sampling [28, 29], where the volume is inferred based on the proportion of sampled points within the polytope. However, these methods are not suitable for computing gradients, as they do not provide a differentiable formulation of the volume. Moreover, maintaining accurate estimates typically requires a prohibitively large number of samples, resulting in significant computational overhead. To address these limitations, we derive analytical expressions for both the volume and its gradient, which admit a predictable polynomial computational complexity (see details in Section IV-D). For an arbitrary base set parameter $\mathbf{h}_{0}$, the volume of the base set defined in (5) can be expressed as
@@
IV-D Computational Complexity and Communication Cost
We now characterize the per-iteration computational cost of the proposed algorithm. In each iteration, the computational workload is divided between the aggregator and all DSRs:
  1. 1.
    Aggregator side. The aggregator performs the projection step (14), and estimates the base-set volume and its gradient. The projection problem is a convex problem with $O(T)$ variables and constraints, solvable in polynomial time using a primal–dual interior-point method. Following Appendices C and D, the $T$-dimensional volume integral (20) and its gradient are computed through a recursive sequence of one-dimensional Riemann sums. Specifically, at each dimension $t$, the state space is discretized into $K_{t}$ grid points, and each grid point aggregates over $K_{t-1}$ points from the previous step, resulting in a total runtime of $O(TK_{\max}^{2})$ with $K_{\max}=\max_{t\in\mathcal{T}}K_{t}$. Hence, the overall per-iteration complexity on the aggregator side is $O(TK_{\max}^{2})+\mathrm{poly}(T)$.
  2. 2.
    DSR side. Each DSR solves its local set-containment problem (12), which is a linear program with $O(T^{2})$ variables and constraints and can be solved in polynomial time with an interior-point method. Since all DSRs operate in parallel, the per-iteration wall-time remains roughly constant with $N$, with per-agent complexity $\mathrm{poly}(T^{2})$.
We then analyze the per-iteration communication cost of the proposed algorithm. In each iteration, the aggregator broadcasts the current base-set parameter $\mathbf{h}_{0}$ to all DSRs, incurring a downlink cost of $O(T)$ scalars that is independent of $N$. Each DSR then returns the optimal affine transformation $\bm{\Gamma}_{i}$ and the associated gradient information. The uplink communication cost per DSR scales as $O(\mathrm{dim}(\bm{\Gamma}_{i}))$, which is $O(T^{2})$ in the worst case, and the total uplink cost therefore scales linearly with $N$. As only aggregated quantities are required and all local computations are parallelized, the overall communication overhead remains lightweight and scalable, while preserving user privacy by design.
IV-E Applicability for Heterogeneous DSRs
While the above pipeline is formulated for resources whose flexibility sets share a common structure, where variability arises primarily from the right-hand-side parameters $\mathbf{h}_{i}$ and all left-hand-side matrices $\mathbf{H}$ are identical, the proposed approach is inherently general and can accommodate a broader range of DSRs. In particular, for devices such as TCLs whose feasible regions depend on both the left-hand and right-hand parameters [7], the same optimization pipeline remains applicable.
Specifically, similar to (3), the flexibility set of TCL $i$ has a left-hand matrix, denoted as $\mathbf{H}_{i}^{\mathrm{TCL}}$, which is fully determined by the dissipation rate $a_{i}$. It can be shown, using Neumann series [30], that $\mathbf{H}_{i}^{\mathrm{TCL}}$ is polynomial in $a_{i}$. The corresponding right-hand-side vector, denoted by $\mathbf{h}_{i}^{\mathrm{TCL}}$, incorporates the modified state and input limits consistent with the parameter $a_{i}$. Thus, an augmented base-set parameter $a_{0},\mathbf{h}_{0}^{\mathrm{TCL}}]$ can be treated as the variable to be optimized. The objective function and its gradient retain structures similar to those in (15) and (17), with corresponding gradient components described in Section IV-C. These quantities can be derived analogously, allowing the optimization to be performed within the same federated base-set framework without modification.
Finally, the proposed framework naturally extends to heterogeneous populations of DSRs, including EVs, TCLs, and other flexible loads, as it does not require identical constraint matrices across participants. Each DSR is modeled by its own polyhedral flexibility set with potentially distinct left-hand and right-hand parameters. Such heterogeneity is handled locally through the set-containment problems, while the aggregator optimizes a shared base set to approximate the collection of individual flexibility sets via affine transformations. When heterogeneity is sufficiently pronounced that a single base set becomes overly restrictive, the framework admits a multi–base-set extension, in which DSRs are grouped by similarity and each group is assigned its own optimized base set. Aggregation then proceeds independently within each group, and the resulting group-level flexibility sets can be either directly used by the power system operator or combined at a higher level, while preserving the same optimization pipeline and privacy properties.
Remark 3 (Applicability for non-linear and discrete dynamics).
The proposed framework can be extended beyond settings where individual DSR models are linear or continuous at the device level. For DSRs with nonlinear dynamics, convex polyhedral inner approximations of the resulting nonconvex but compact and continuous sets have been established in the literature [31, 32], and can be directly used within the framework. For DSRs with discrete control actions (e.g., ON/OFF switching in TCLs), continuous convex flexibility sets can be obtained by modeling flexibility at an appropriate temporal or population scale, where binary actions admit meaningful continuous representations (e.g., duty cycles or population-level averages) [7, 33]. As long as a DSR’s flexibility can be represented by a convex polytope in H-representation, our framework remains applicable.
@@ VI-A1 Parameter settings
We consider a finite time horizon with each time interval being 1 hour ($\Delta t=1$), and an EV fleet consisting of 50 vehicles ($N=50$). The parameter intervals of EVs are summarized in Table I (the parameters of a specific EV are uniformly sampled from these intervals). The energy demand for EV $i$ is computed by first determining its maximum feasible charge, given by $u_{i}^{\mathrm{max}}(t_{i}^{\mathrm{d}}-t_{i}^{\mathrm{p}})$, and then scaling it by a random number drawn uniformly from $[0.2,0.5]$. The parameter $\epsilon$ in constraint (14b) is set to $1\times 10^{-5}$, which is sufficiently small to ensure strict feasibility without affecting the geometry of the base set.
@@
VI-A2 Initialization and step-size selection
The gradient-based algorithm in Section IV is initialized with an all-one vector, providing a feasible starting point [^5] This initialization guarantees feasibility of (13b), since the origin $\mathbf{u}=\mathbf{0}$ strictly satisfies $\mathbf{H}\mathbf{u}\leq\mathbf{1}$, ensuring that the base set has a nonempty interior.. The step-size selection follows a standard backtracking line search method [26] to ensure monotonic ascent of the objective function.
VI-A3 Benchmarks
@@
In addition to the baseline AVG, we also include a centralized full-information benchmark in Section VI-C, where the aggregator directly solves the original optimization problems with access to all individual flexibility sets. This serves as a best-case performance reference rather than a deployable baseline.
@@ VI-B Volume of Aggregate Flexibility Set
To further evaluate the efficiency of the proposed federated base-set optimization framework, we investigate its convergence behavior and scalability with respect to the number of participating agents. The convergence performance is quantified using the normalized suboptimality defined as
$e=\frac{J(\mathbf{h}_{0}^{\mathrm{opt}})-J\big(\mathbf{h}_{0}^{(k)}\big)}{J(\mathbf{h}_{0}^{\mathrm{opt}})},$ (26)
where $J\big(\mathbf{h}_{0}^{(k)}\big)$ denotes the objective value at iteration $k$. This metric measures the relative distance from the current iterate to the optimal solution. Fig. 4 illustrates the convergence trajectories of $e$ for different numbers of EVs ($N=20,30,50$). Only the first ten iterations are displayed, as the objective values in all cases stabilize within this range, indicating convergence or near-convergence. The average per-iteration computation time at the aggregator side is 42.7 s, 37.4 s, and 44.1 s for $N=20$, 3030, and $50$, respectively, while the corresponding average per-iteration computation time at the DSR side is 9.8 s, 11.2 s, and 7.6 s. These results indicate that both the convergence speed and per-iteration computational cost exhibit limited sensitivity to different values of $N$, confirming that the proposed algorithm scales efficiently with the number of participating EVs.
Figure 4: Convergence trajectories of the normalized error $e$ for different numbers of EVs $N=20,30,50$.
@@ VI-C2 Electricity cost minimization
$\Pi^{\star}=\min_{\mathbf{u}_{i},\forall i\in\mathcal{N}}\ \bigg(\!\bm{\pi}^{\top}\sum_{i\in\mathcal{N}}\mathbf{u}_{i}\!\bigg)\Delta t\ \ \text{s.t. }\ \mathbf{u}_{i}\in\mathcal{U}_{i},\,\forall i\in\mathcal{N}.$ (30)
Fig. 7 illustrates i) the relative electricity cost gaps of the AVG and Proposed methods against the best-case benchmark, computed as $\big(\Pi(\mathbf{h}_{0}^{\mathrm{avg}})-\Pi^{\star}\big)/{\Pi^{\star}}$ and $\big(\Pi(\mathbf{h}_{0}^{\mathrm{opt}})-\Pi^{\star}\big)/{\Pi^{\star}}$, respectively, and ii) the electricity cost improvement of the Proposed method over the AVG method, measured by the relative reduction $\big(\Pi(\mathbf{h}_{0}^{\mathrm{avg}})-\Pi(\mathbf{h}_{0}^{\mathrm{opt}})\big)/{\Pi(\mathbf{h}_{0}^{\mathrm{avg}})}$. While both methods incur some level of optimality gaps compared to the best-case benchmark, our Proposed method consistently demonstrates superior performance. Employing the optimized base-set parameter $\mathbf{h}_{0}^{\mathrm{opt}}$ markedly narrows the gap. As highlighted by the third boxplot, this translates into a substantial relative cost savings, with our approach consistently outperforming the AVG benchmark by a significant margin. These results provide strong evidence that a more accurate flexibility characterization, as enabled by our method, directly translates into economic value.
@@ Appendix B Gradient Derivation for Set Containment Problem
$\nabla_{\bm{\gamma}}\mathcal{L}=\mathbf{H}^{\top}\bm{\mu}^{\star}=\mathbf{0},\quad\!\nabla_{\bm{\Gamma}}\mathcal{L}=-\mathbf{I}_{T}-\mathbf{H}^{\top}\mathbf{Y}^{\star}=\mathbf{0},$ (36a)
$\nabla_{\bm{\Lambda}}\mathcal{L}=\mathbf{Y}^{\star}\mathbf{H}^{\top}+\bm{\mu}^{\star}{\mathbf{h}}_{0}^{\top}-\bm{\nu}^{\star}=\mathbf{0},$ (36b)
$\bm{\Lambda}^{\star}\geq\bm{0},\quad\bm{\Lambda}^{\star}\mathbf{H}-\mathbf{H}\bm{\Gamma}^{\star}=\mathbf{0},$ (36c)
$\mathbf{g}^{\star}=\bm{\Lambda}^{\star}{\mathbf{h}}_{0}+\mathbf{H}\bm{\gamma}^{\star}-\mathbf{h}\leq\mathbf{0},$ (36d)
$\bm{\mu}^{\star}\geq\mathbf{0},\quad\bm{\nu}^{\star}\geq\mathbf{0},\quad\bm{\mu}^{\star}\circ\mathbf{g}^{\star}=\mathbf{0},\quad\bm{\nu}^{\star}\circ\bm{\Lambda}^{\star}=\mathbf{0},$ (36e)